php?id-7422845 ng Jump to... Rob n Stahe University-ECON 2100 Microeconomics- Spring19-KHEIRANDISH Activities and Due Dates ) Appendax A Problem Set < ⓘ Assignment Score: Hint Question 2 of 12 For the straight line defined by the points (3,53) and (5,87), determine the slope (m) and y-intercept (b). Do not round the answers ← Backspace CV B N t shit
Answer #1

Slope can be given as m= Y2-Y1/X2-X1

M= 87-53/5-3

M= 34/2

M=17

For intercept,

We know the eqn of line can be given as

Y = MX + C

53 = 17*3 + C

C = 53-51= 2